PIECRWISE PROBLEMS
Mark started selling snacks in the nearby school.in one day he spend 200 for rent and 25 for each snack item he prepares. His expenses in a single day can expressed as function C(x)=+200,where x is the number of item and C(x) is his daily expenses in pesos.how much are his expenses if he prepare 100 snack item?150 snack item?

Answer by Theo(13342)

he spends 200 a day for rent.
his expenses for making each snack are .25 per snack.

his cost function is:

c(x) = 200 + .25 * x

c(x) is the total vost for the day.

if he makes 100 snacks, c(x) = 200 + .25 * 100 = 200 + 25 = 225.

if he makes 150 snacks, c(x) = 200 + .25 * 150 = 237.5.

you can graph the function.

it looks like this.